complete the square.
y=x^2-4x+2

Answer by stanbon(75887)

complete the square.
y=x^2-4x+2
-------------
x^2-4x = -y-2
x^2-4x+4 = -y-2+4
(x-2)^2 = -(y-2)
=====================
Cheers,
Stan H.